It is necessary to add iodide ions to precipitate the lead(II) ions from 250 mL of 0.03 M Pb(NO3)2(aq). What minimum iodide ion concentration is required for the onset of PbI2 precipitation? The solubility product of PbI2 is 1.4 × 10−8. (Answer in units of mol/L.)


What mass of KI must be added for PbI2 to form? Answer in units of g.


1
Expert's answer
2021-03-11T08:48:16-0500

PBI2 "\leftrightarrow" Pb2+ _ 2I-

250 mL of0.03M

[Pb2+] [I-]2 =Ksp

0.03 (x2) = 1.4(10-8)

x= 6.83 (10-4)M

moles of KI = 6.84(10-4)M (250 mL)

= 0.17 mmol

mass of KI required = 0.17 (166) = 28.22 mg

28.22mg/1000 = 0.02822g